### v2

# MIGHTY: Hermite Spline-based Efficient Trajectory Planning

Abstract
Hard-constraint trajectory planners often rely on commercial solvers and demand substantial computational resources. Existing soft-constraint methods achieve faster computation, but either (1) decouple spatial and temporal optimization or (2) restrict the search space. To overcome these limitations, we introduce MIGHTY, a Hermite spline-based planner that performs spatiotemporal optimization while fully leveraging the continuous search space of a spline. In simulation, MIGHTY achieves a 9.3% reduction in computation time and a 13.1% reduction in travel time over state-of-the-art baselines, with a 100% success rate. In hardware, MIGHTY completes multiple high-speed flights up to 6.7 m/s in a cluttered static environment and long-duration flights with dynamically added obstacles.
Supplementary Material
I Introduction
Trajectory planning for autonomous navigation has been extensively studied, with a wide variety of parameterizations and formulations [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus, zhou2019robust, zhou2021ego-planner, zhou2022swarm, zhou2021raptor, tordesillas2022faster, richter2016polynomial, levy2024stitcher, wang2022geometrically, ren2025super]. For real-world deployment, fast trajectory generation is critical to ensure agile and safe flight in cluttered and dynamic environments. Approaches that use hard constraints [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus] explicitly enforce safety and physical feasibility but often rely on computationally intensive solvers, making them difficult to deploy for high-frequency replanning on onboard processors. Soft-constraint planners such as EGO-Planner [zhou2021ego-planner], RAPTOR [zhou2021raptor], and SUPER [ren2025super] achieve faster convergence and lower computational cost. Some jointly optimize path geometry and timing within a single nonlinear formulation [wang2022geometrically, ren2025super], while others separate path generation and time allocation into two stages [richter2016polynomial, zhou2021raptor]. In general, increasing the number of decision variables, such as control points, segment durations, or derivative values, can improve performance but also enlarges the optimization problem and increases computation time. To balance efficiency and performance, recent work has explored alternative trajectory representations and optimization strategies, and building on these ideas, the following introduces our Hermite spline–based planner, MIGHTY.
TABLE I: State-of-the-art Unconstrained UAV Trajectory Planners
Method Representation Variables Time Opt. Search Space Local Control
Polynomial [richter2016polynomial] (2016) 9th-order polynomial endpoint derivs, time decoupled limited polynomial+time space yes
Robust [zhou2019robust] (2019) non-uniform cubic B-spline B-spline cntrl pts, time decoupled limited polynomial+time space yes
EGO-Planner [zhou2021ego-planner] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
RAPTOR [zhou2021raptor] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
MINCO class [wang2022geometrically, zhou2022swarm, ren2025super] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO manifold) no
MIGHTY (proposed) Hermite Hermite cntrl pts, time. joint full polynomial+time space yes
I-A Review of Unconstrained Frameworks
Differential flatness [FLIESS01061995] enables efficient optimization of quadrotor trajectories by parameterizing the flat output ($x$, $y$, $z$, yaw) and their derivatives. Given fixed timestamps at trajectory waypoints, minimum control effort (e.g., jerk/snap) problems using piecewise polynomial representation can be formulated as quadratic programs (QPs), which are convex and can be solved efficiently [mellinger2011minimum]. However, when segment durations are included as optimization variables (time allocation), the resulting joint problem becomes substantially more complex, yielding a nonlinear program (NLP) that is often ill-conditioned because of the coupling between spatial and temporal variables.
One common strategy is to decouple the problem into two stages and relax hard constraints to adopt unconstrained formulations. For example, Richter et al. [richter2016polynomial] apply piecewise 9th-order polynomials on the flat outputs with endpoint derivatives as decision variables. The trajectory optimization follows a two-stage process: an unconstrained QP minimizes snap to ensure smoothness, followed by a gradient-based time allocation step that adjusts the total trajectory duration. Collision avoidance is addressed heuristically by inserting mid-segment waypoints whenever a segment intersects obstacles, which triggers repeated QP solves until a collision-free solution is found. This approach is extended in [7353622] by improving numerical stability and incorporating additional soft state constraint costs to further ensure trajectory feasibility. Similarly, the work in [7759784] adapts this framework to achieve local real-time replanning, using endpoint derivatives as optimization variables, which minimizes snap and collisions using a potential cost function based on Euclidean signed distance field (ESDF).
Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector. B-splines offer inherent smoothness, local control, and the convex hull property, which enables efficient collision checking and enforces dynamic feasibility [9147300, zhou2019robust, zhou2021ego-planner, zhou2021raptor]. Zhou et al. [zhou2019robust] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths. The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration. Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility. EGO-Planner [zhou2021ego-planner] and RAPTOR [zhou2021raptor] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points. EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [liu1989lbfgs] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness. RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
To further improve efficiency, alternative representations of polynomial variables have been explored. Wang et al. [wang2021glst] employ an explicit diffeomorphism between coefficients and boundary derivatives, with a linear-complexity scheme for gradient evaluation towards coefficients. Built on this, MINCO [wang2022geometrically] parametrizes a segmented trajectory as piecewise polynomials, where coefficients are analytically determined by intermediate waypoints and segment durations, which are jointly optimized. A linear-complexity solver then constructs the minimum-control trajectory that minimizes the integrated control effort (e.g., jerk/snap) subject to flatness boundary conditions. Spatiotemporal deformation leverages smooth diffeomorphisms and convex constraint-elimination maps to enforce arbitrary collision and state-input constraints. However, although MINCO provides an elegant and computationally efficient formulation, since it directly maps intermediate waypoints and segment durations to polynomial coefficients using the optimality conditions of the optimal control-effort problem [wang2022geometrically], the search is, by construction, restricted to a low-dimensional manifold, potentially limiting flexibility when additional constraints or objectives are introduced. Additionally, because MINCO determines all polynomial coefficients by solving a single banded linear system subject to high-order continuity and optimality conditions, any change to one intermediate waypoint or segment duration generally changes the entire solution. In other words, MINCO does not have strict local control, meaning a small change to one waypoint or duration can propagate to many segments, complicating per-segment constraint handling for local changes. EGO-Swarm2 [zhou2022swarm] and SUPER [ren2025super] both adopt MINCO to compactly parameterize trajectories, with SUPER further optimizing exploratory and backup trajectories in each replanning cycle to balance aggressiveness and guaranteed safety. However, the intrinsic coupling within the MINCO formulation remains a limitation. MIGHTY aims to address this by optimizing over a higher-dimensional representation that preserves local control and improves flexibility for handling complex constraints.
I-B MIGHTY Contributions
Summarizing the discussion above, direct unconstrained higher-order polynomial representations [richter2016polynomial] are typically too slow for real-time planning. Uniform or Non-uniform splines [zhou2019robust, zhou2021ego-planner, zhou2021raptor] can generate trajectories faster compared to spatiotemporal joint optimization; however, they can generate sub-optimal trajectories due to decoupled time adjustment. Waypoint-and-duration-based parameterizations [wang2022geometrically, zhou2022swarm, ren2025super] are restricted to a low-dimensional manifold, and a small change in one waypoint propagates globally through all polynomial coefficients. This is undesirable for real-time planning because it potentially complicates constraint/cost handling and slows down convergence.
To address the limitations of existing soft-constraint planners, MIGHTY uses a Hermite-spline representation and an unconstrained nonlinear optimizer that searches directly over spatial waypoints, endpoint derivatives, and segment durations. A Hermite spline representation also guarantees continuity without global coefficient coupling. As such, the key contributions of the paper are:
  1. 1.
    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, fully leveraging the continuous search space of Hermite splines and supporting local control of trajectory shape. The approach, MIGHTY, achieves fast solve times and superior performance.
  2. 2.
    Comprehensive Simulation Study: We benchmark in both simple and complex scenes, comparing against state-of-the-art methods in static environments and validating safe behavior in dynamic environments. MIGHTY achieves $9.3\text{\,}\%$ reduction in computation time and $13.1\text{\,}\%$ reduction in travel time in trajectory representation benchmarking, and it achieves the shortest travel time and path length while maintaining the $100\text{\,}\%$ success rate.
  3. 3.
    Extensive Hardware Experiments: Long-duration flights, high-speed flights, and dynamic obstacle avoidance scenarios using a LiDAR-based perception and localization system. MIGHTY achieves collision-free long-duration flights and a top speed of $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ in high-speed flight experiments.
II Hermite Spline
This section introduces the Hermite-spline notation used in MIGHTY’s formulation. MIGHTY uses a degree-$d$ Hermite spline with $M$ segments. For odd degree $d=2k+1$, each segment is parameterized by the knot values of the position and its derivatives up to order $k$. For example, a quintic Hermite spline ($d=5$, $k=2$) specifies position, velocity, and acceleration. Across knots, a Hermite spline automatically guarantees $\mathcal{C}^{k}$ continuity by sharing the knot values. Although Hermite splines apply in any dimension, we work in $\mathbb{R}^{3}$ for quadrotor planning and use $d=5$ in this paper. For each segment $s=0,\dots,M-1$, we define a normalized time variable $\tau_{s}\in[0,1]$ on segment $s$ as $\tau_{s}=(t-t_{s})/T_{s}\,\in[0,1].$ Then each segment can be denoted as: $\mathbf{x}(t)=\mathbf{x}_{s}(\tau_{s}),\,t\in\bigl[t_{s},\;t_{s+1}\bigr],$ where $t_{s}:=\sum_{r=0}^{s-1}T_{r}$, and $T_{s}$ is the duration of segment $s$. Then we write
$\quad\mathbf{x}_{s}(\tau_{s})=\sum_{k=0}^{5}h_{k}(\tau_{s})\,\mathbf{H}_{s,k},$ (1)
where $\mathbf{H}_{s,0}=\mathbf{p}_{s},\mathbf{H}_{s,1}=T_{s}\,\mathbf{v}_{s},\mathbf{H}_{s,2}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s},\mathbf{H}_{s,3}=\mathbf{p}_{s+1},\ \mathbf{H}_{s,4}=T_{s}\,\mathbf{v}_{s+1},\ \mathbf{H}_{s,5}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s+1}.$ where $h_{k}(\tau_{s})$ are the standard quintic Hermite basis functions. The $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}\in\mathbb{R}^{3}\;(i=0,\dots,M)$ and $T_{s}>0\;(s=0,\dots,M-1)$ are the interior knot positions, velocities, accelerations, and segment durations, respectively.
III MIGHTY
III-A Decision Variables
As shown in Section II, the knot states $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}(i=0,\dots,M)$ and segment durations $T_{s}(s=0,\dots,M-1)$ fully determine the trajectory. Thus we optimize the interior positions, velocities, accelerations, and per-segment durations:
$\mathbf{z}=\bigl[\,\mathbf{p}_{1}^{\top},\mathbf{v}_{1}^{\top},\mathbf{a}_{1}^{\top},\ldots,\mathbf{p}_{M-1}^{\top},\mathbf{v}_{M-1}^{\top},\mathbf{a}_{M-1}^{\top},T_{0},\ldots,T_{M-1}\bigr].^{\top}$
with fixed boundary states $(\mathbf{p}_{0},\mathbf{v}_{0},\mathbf{a}_{0})$ and $(\mathbf{p}_{M},\mathbf{v}_{M},\mathbf{a}_{M})$.
To keep the optimization unconstrained while enforcing certain requirements (e.g., $T_{s}>0$), we apply diffeomorphism, such as $T_{s}\;=\;\phi(\sigma_{s})$ with $\phi(\sigma)=e^{\sigma},\ \sigma_{s}\in\mathbb{R}.$ We also find that optimizing raw derivative knots $(\mathbf{v}_{i},\mathbf{a}_{i})$ can be numerically unstable. Section III-D introduces scalings that improve stability without changing the optimum. For clarity, the formulas below are written in terms of the original variables $(\mathbf{p}_{s},\mathbf{v}_{s},\mathbf{a}_{s},T_{s})$; applying a diffeomorphism only inserts chain-rule factors in the gradients.
III-B Representations
Given the quintic Hermite segment in Eq. (1), a smooth cost can be expressed in terms of the raw Hermite variables $(\mathbf{p}_{s},\mathbf{v}_{s},\mathbf{a}_{s},\mathbf{p}_{s+1},\mathbf{v}_{s+1},\mathbf{a}_{s+1},T_{s})$. However, evaluating costs and gradients directly in Hermite form can be computationally expensive, and other representations, such as Bézier curve, can be more efficient.
For instance, a Bézier curve allows efficient evaluation of the trajectory and its derivatives at arbitrary sample points. A degree-$n$ Bézier segment is written as $\mathbf{x}(u)=\sum_{i=0}^{n}B_{i}^{n}(u)\,\mathbf{c}_{i},u\in[0,1],$ where the vectors $\mathbf{c}_{i}$ are the Bézier control points, and the Bernstein basis polynomials are $B_{i}^{n}(u)\;=\;\binom{n}{i}\,u^{\,i}\,(1-u)^{\,n-i},i=0,\dots,n.$ Because evaluation is just a weighted sum of control points with weights $B_{i}^{n}(u)$, sampling reduces to a few dot products with small, reusable tables. Then each sample is a small dot product: positions use the $B_{i}^{5}$ table, and derivatives use the degree-reduced Bernstein weights. Since these tables are tiny and shared across iterations, and each dot product has only $6,5,4,3$ terms (for degree $5$ down to $2$), this is faster and numerically steadier than recomputing higher-order Hermite polynomials and their derivatives at every sample.
Although Bézier representation has these computational advantages, note that Bézier control points do not enforce cross-segment continuity by themselves. On the other hand, Hermite variables guarantee $\mathcal{C}^{k}$ continuity across segments by construction. Therefore, it is often advantageous to optimize in the Hermite parameterization while evaluating cost terms in the Bézier basis to exploit its computational benefits. In our simulations and hardware experiments, we optimize over Hermite and evaluate costs in Bézier; however, in general, MIGHTY does not require Bézier for cost evaluation.
III-C Objective and Closed-Form Gradient
To make the formulation basis-independent, we express the objective and its gradients with respect to a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.). As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
$\mathbf{c}_{s,0}$ $=\mathbf{p}_{s},\ \mathbf{c}_{s,1}=\mathbf{p}_{s}+\tfrac{T_{s}}{5}\,\mathbf{v}_{s},\ \mathbf{c}_{s,2}=\mathbf{p}_{s}+\tfrac{2T_{s}}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}^{2}}{20}\,\mathbf{a}_{s},$ (2)
$\mathbf{c}_{s,3}$ $=\mathbf{p}_{s+1}-\tfrac{2T_{s}}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}^{2}}{20}\,\mathbf{a}_{s+1},\ \mathbf{c}_{s,4}=\mathbf{p}_{s+1}-\tfrac{T_{s}}{5}\,\mathbf{v}_{s+1},$
$\mathbf{c}_{s,5}$ $=\mathbf{p}_{s+1}.$
Let $\mathbf{c}_{s}=[\mathbf{c}_{s,0}^{\top}\!\cdots\!\mathbf{c}_{s,5}^{\top}]^{\top}$ and $\mathbf{y}_{s}=[\mathbf{p}_{s}^{\top},\mathbf{v}_{s}^{\top},\mathbf{a}_{s}^{\top},\mathbf{p}_{s+1}^{\top},\mathbf{v}_{s+1}^{\top},\mathbf{a}_{s+1}^{\top}]^{\top}$, so $\mathbf{c}_{s}=C(T_{s})\,\mathbf{y}_{s}$. Let the total objective be $J(\mathbf{z})\;=\;\sum_{s=0}^{M-1}J_{s}\bigl(\mathbf{c}_{s}\bigr).$ To obtain $\partial J/\partial\mathbf{z}$, we first compute the derivatives w.r.t $\mathbf{y}_{s}$ and $T_{s}$, then apply the chain rule back to $\mathbf{z}$. Denote $\mathbf{g}_{\mathbf{c}_{s}}\equiv\partial J_{s}/\partial\mathbf{c}_{s}$ and the explicit time derivative $\bigl.\partial J_{s}/\partial T_{s}|_{\text{explicit}}.$ Since ${\partial\mathbf{c}_{s}}/{\partial\mathbf{y}_{s}}=C(T_{s})$, the chain rule yields
$\frac{\partial J_{s}}{\partial\mathbf{y}_{s}}\;$ $=\;C(T_{s})^{\top}\,\mathbf{g}_{\mathbf{c}_{s}},$ (3)
$\frac{\partial J_{s}}{\partial T_{s}}\;$ $=\;\Bigl.\frac{\partial J_{s}}{\partial T_{s}}\Bigr|_{\text{explicit}}\;+\;\mathbf{g}_{\mathbf{c}_{s}}^{\top}\frac{\partial\mathbf{c}_{s}}{\partial T_{s}}.$ (4)
Note that the explicit time derivative $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}$ comes from factors like $T_{s}^{5}$ in smoothness integrals. Using Eq. (2), the entries of $C(T_{s})$ are simple constants, so Eq. (3) expands to the following closed-form expressions:
$\nabla_{\mathbf{p}_{s}}J_{s}$ $=\mathbf{g}_{\mathbf{c}_{s,0}}+\mathbf{g}_{\mathbf{c}_{s,1}}+\mathbf{g}_{\mathbf{c}_{s,2}},\nabla_{\mathbf{v}_{s}}J_{s}=\tfrac{T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,1}}+\tfrac{2T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,2}},$
$\nabla_{\mathbf{a}_{s}}J_{s}$ $=\tfrac{T_{s}^{2}}{20}\,\mathbf{g}_{\mathbf{c}_{s,2}},\nabla_{\mathbf{p}_{s+1}}J_{s}=\mathbf{g}_{\mathbf{c}_{s,3}}+\mathbf{g}_{\mathbf{c}_{s,4}}+\mathbf{g}_{\mathbf{c}_{s,5}},$
$\nabla_{\mathbf{v}_{s+1}}J_{s}$ $=-\tfrac{2T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,3}}-\tfrac{T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,4}},\nabla_{\mathbf{a}_{s+1}}J_{s}=\tfrac{T_{s}^{2}}{20}\,\mathbf{g}_{\mathbf{c}_{s,3}}.$
And $\partial\mathbf{c}_{s}/\partial T_{s}$ used in the coefficient path in Eq. (4) can be computed in closed form as
$\frac{\mathbf{c}_{s,0}}{T_{s}}$ $=0,\quad\frac{\mathbf{c}_{s,1}}{T_{s}}=\tfrac{1}{5}\,\mathbf{v}_{s},\quad\frac{\mathbf{c}_{s,2}}{T_{s}}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
$\frac{\mathbf{c}_{s,3}}{T_{s}}$ $=-\tfrac{2}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s+1},\quad\frac{\mathbf{c}_{s,4}}{T_{s}}=-\tfrac{1}{5}\,\mathbf{v}_{s+1},\quad\frac{\mathbf{c}_{s,5}}{T_{s}}=0.$
Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Section IV.
III-C1 Cost on Control Points
We first consider cost terms that are expressed directly in terms of the coefficients/control points, such as integrated squared jerk. For a quintic Bézier segment $s$ with control points $\mathbf{c}_{s,0},\ldots,\mathbf{c}_{s,5}\in\mathbb{R}^{3}$, obtained from the Hermite endpoint states via Eq. (2), the third forward differences are: $\Delta_{s,m}=\mathbf{c}_{s,m+3}-3\,\mathbf{c}_{s,m+2}+3\,\mathbf{c}_{s,m+1}-\mathbf{c}_{s,m},\ m\in\{0,1,2\}.$ Then, the integrated squared jerk on segment $s$ is
$J_{\text{smooth},s}\!=\!\int_{0}^{T_{s}}\!\!\|\mathbf{j}(t)\|^{2}\,dt\!=\!C_{s}\int_{0}^{1}\!\Big\|\sum_{m=0}^{2}\Delta_{s,m}\,B_{m}^{2}(\tau_{s})\Big\|^{2}d\tau_{s},$
where $C_{s}=3600\,T_{s}^{-5}$. Define the (scalar) Bernstein-quadratic Gram matrix $G\in\mathbb{R}^{3\times 3}$ by $G_{mn}\;\triangleq\;\int_{0}^{1}B_{m}^{2}(\tau_{s})\,B_{n}^{2}(\tau_{s})\,d\tau_{s}\quad(m,n=0,1,2),$ so that, with $\Delta_{s}\triangleq$ then $\Delta_{s,0}^{\top},\Delta_{s,1}^{\top},\Delta_{s,2}^{\top}]^{\top}\in\mathbb{R}^{9}$,
$J_{\text{smooth}}\;=\;\sum_{s=0}^{M-1}C_{s}\;\Delta_{s}^{\top}\,Q\,\Delta_{s},\qquad Q\;\triangleq\;G\otimes I_{3}.$ (6)
(Explicitly, $G$ is symmetric positive definite with constants $G_{00}=G_{22}=\tfrac{1}{5},\;G_{11}=\tfrac{2}{15},\;G_{01}=G_{12}=\tfrac{1}{10},\;G_{02}=\tfrac{1}{30}$.)
Let $\lambda_{\min},\lambda_{\max}>0$ be the smallest and largest eigenvalues of $G$. Since $Q=G\otimes I_{3}$, the eigenvalues of $Q$ are the eigenvalues of $G$, each repeated three times. Hence $\lambda_{\min}\,\|\Delta_{s}\|^{2}\;\leq\;\Delta_{s}^{\top}Q\Delta_{s}\;\leq\;\lambda_{\max}\,\|\Delta_{s}\|^{2}.$ Therefore a computationally cheap surrogate is the diagonally weighted form $\tilde{J}_{\text{smooth}}\;=\;\sum_{s=0}^{M-1}C_{s}\sum_{m=0}^{2}w_{m}\,\|\Delta_{s,m}\|^{2},$ with fixed positive weights $w_{m}$, which is equivalent to the exact cost up to a constant factor bounded by $\lambda_{\min}$ and $\lambda_{\max}$.
Now we derive $\partial J_{\text{smooth}}/\partial\mathbf{c}$: Each $\Delta_{s,m}$ is an affine combination of four consecutive control points with coefficients $\alpha_{j,m}\in\{-1,3,-3,1\}$ at indices $j=m,\ldots,m+3$. Hence, $\partial J_{\text{smooth}}/{\partial\Delta_{s,m}}=2\,C_{s}\,\Delta_{s,m}$ $\partial J_{\text{smooth}}/{\partial\mathbf{c}_{s,j}}=2\,C_{s}\sum_{m=0}^{2}\alpha_{j,m}\,\Delta_{s,m}.$ Stacking, we obtain
$\mathbf{g}_{\mathbf{c}_{s}}=\frac{\partial J_{\text{smooth}}}{\partial\mathbf{c}_{s}}=\big[(\frac{\partial J_{\text{smooth}}}{\partial\mathbf{c}_{s,0}})^{\top}\cdots(\frac{\partial J_{\text{smooth}}}{\partial\mathbf{c}_{s,5}})^{\top}\big]^{\top}.$ (7)
We can now derive $\partial J_{\text{smooth}}/\partial T_{s}$. Eq. (4) has two contributions. Since $C_{s}=3600\,T_{s}^{-5}$, the first contribution is
$\Bigl.\frac{\partial J_{s}}{\partial T_{s}}\Bigr|_{\text{explicit}}=-\frac{5}{T_{s}}\,J_{{\mathrm{smooth}},s}.$ (8)
Using Eqs. (5) and (7), we derive the second term of Eq. (4) and obtain a closed-form expression.
III-C2 Cost on Sampled States
Now we consider cost terms that are evaluated on sampled states of the trajectory. Fix a set of sample points $\{\tau_{s,j}\}_{j=0}^{\kappa_{s}}\subset[0,1]$ for segment $s$. Define a generic per-sample cost $\ell:\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}_{>0}\to\mathbb{R}$, $\ell(\mathbf{x},\mathbf{v},\mathbf{a},\mathbf{j};T)$. The sampled cost on segment $s$ is
$J_{\text{samp},s}\;=\;\sum_{j=0}^{\kappa_{s}}\ell\!\bigl(\mathbf{x}_{s}(\tau_{s,j}),\mathbf{v}_{s}(\tau_{s,j}),\mathbf{a}_{s}(\tau_{s,j}),\mathbf{j}_{s}(\tau_{s,j});\,T_{s}\bigr).$
Let the sample Jacobians at $\tau_{s,j}$ be $\ell_{\mathbf{x}}=\partial\ell/\partial\mathbf{x}$, $\ell_{\mathbf{v}}=\partial\ell/\partial\mathbf{v}$, $\ell_{\mathbf{a}}=\partial\ell/\partial\mathbf{a}$, $\ell_{\mathbf{j}}=\partial\ell/\partial\mathbf{j}$. Typically $\ell$ has no explicit $T_{s}$ (as above), so $T_{s}$ appears only via time-scaling and quadrature weights.  [^1] If an explicit time term is desired, then $\ell=\ell(\mathbf{x},\mathbf{v},\mathbf{a},\mathbf{j};T_{s})$ and an extra $\partial\ell/\partial T_{s}$ term would be added when differentiating w.r.t. $T_{s}$. Then, holding $T_{s}$ fixed, the gradient w.r.t. the $k$th control point is
$\mathbf{g}_{c_{s},k}$ $=\sum_{j=0}^{\kappa_{s}}\,\Bigl[B_{k}^{5}\,\ell_{\mathbf{x}}+\tfrac{5}{T_{s}}\bigl(B_{k-1}^{4}-B_{k}^{4}\bigr)\,\ell_{\mathbf{v}}$
$+\tfrac{20}{T_{s}^{2}}\bigl(B_{k-2}^{3}-2B_{k-1}^{3}+B_{k}^{3}\bigr)\,\ell_{\mathbf{a}}$
$+\tfrac{60}{T_{s}^{3}}\bigl(B_{k-3}^{2}-3B_{k-2}^{2}+3B_{k-1}^{2}-B_{k}^{2}\bigr)\,\ell_{\mathbf{j}}\Bigr],$
where we omit the argument of the Bernstein basis polynomials $B(\tau_{s,j})$ for clarity. Note that $B_{r}^{n}(\cdot)=0$ if $r\notin\{0,\dots,n\}$. Stacking over $k=0,\ldots,5$ gives $\mathbf{g}_{\mathbf{c}_{s}}=\partial J_{\text{samp},s}/\partial\mathbf{c}_{s}$, which is then pulled back (chain rule) and accumulated into the Hermite variables via Eq. (3).
[…]
  • (a)
    State-scaling (fix $\mathbf{c}_{s}$). Holding $\mathbf{c}_{s}$ fixed means $\mathbf{x}(\tau)$ is unchanged at each sample $\tau_{s,j}$, but $\frac{\partial\mathbf{v}}{\partial T_{s}}=-\frac{1}{T_{s}}\mathbf{v},\quad\frac{\partial\mathbf{a}}{\partial T_{s}}=-\frac{2}{T_{s}}\mathbf{a},\quad\frac{\partial\mathbf{j}}{\partial T_{s}}=-\frac{3}{T_{s}}\mathbf{j}.$ Thus,
    $\hskip-18.06749pt\Bigl.\frac{\partial J_{\text{samp},s}}{\partial T_{s}}\Bigr|_{\text{states}}=\sum_{j=0}^{\kappa_{s}}\,\Bigl[-\frac{1}{T_{s}}\,\ell_{\mathbf{v}}\mathbf{v}-\frac{2}{T_{s}}\,\ell_{\mathbf{a}}\mathbf{a}-\frac{3}{T_{s}}\,\ell_{\mathbf{j}}\mathbf{j}\Bigr]_{\tau=\tau_{s,j}}.$
  • (b)
    Coefficient. In Hermite parameterizations, the Bézier control points depend on $T_{s}$ (e.g., tangent/curvature control points include factors of $T_{s}$ and $T_{s}^{2}$). Using Eq. (4) gives $\Bigl.\frac{\partial J_{\text{samp},s}}{\partial T_{s}}\Bigr|_{\text{coeff}}=\mathbf{g}_{c_{s}}^{\top}\,\frac{\partial\mathbf{c}_{s}}{\partial T_{s}}.$
Thus, combining (a) & (b) above yields, $\frac{\partial J_{\text{samp},s}}{\partial T_{s}}=\Bigl.\frac{\partial J_{\text{samp},s}}{\partial T_{s}}\Bigr|_{\text{states}}+\Bigl.\frac{\partial J_{\text{samp},s}}{\partial T_{s}}\Bigr|_{\text{coeff}}.$ Note that since generally sampling cost $\ell$ does not depend on $T_{s}$ explicitly, $\partial J_{s}/{\partial T_{s}}|_{\text{explicit}}=0$.
III-D Reparameterizations for Derivative Variables
Each segment with duration $T_{s}$ is parameterized by normalized time $\tau_{s}\in[0,1]$; by the chain rule,
$\mathbf{v}(t)=\frac{\partial\mathbf{x}}{\partial t}=\frac{1}{T_{s}}\frac{\partial\mathbf{x}}{\partial\tau_{s}},\qquad\mathbf{a}(t)=\frac{\partial^{2}\mathbf{x}}{\partial t^{2}}=\frac{1}{T_{s}^{2}}\frac{\partial^{2}\mathbf{x}}{\partial\tau_{s}^{2}}.$
Optimizing raw knot derivatives $\{\mathbf{v}_{i},\mathbf{a}_{i}\}$ together with $\{T_{s}\}$ introduces explicit $1/T_{s}$ and $1/T_{s}^{2}$ factors in the gradients, so short segments have larger influence. We therefore optimize scaled knot derivatives. Define an averaged local time as
$\bar{T}_{i}=\begin{cases}T_{0},&i=0,\\ \frac{T_{i-1}+T_{i}}{2},&i=1,\dots,M-1,\\ T_{M-1},&i=M,\end{cases}$
and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization removes the explicit $1/T_{s}$ factors from the gradients and balances sensitivity across short and long segments. Our ablation study in Section IV-C confirms the expected improvement in performance.
IV Simulation Results
All simulations were run on an AlienWare Aurora R8 with an Intel® CoreTM i9-9900K CPU and 64 GB RAM, using Ubuntu 22.04 LTS and ROS 2 Humble.
IV-A Representation Benchmarking: Simple Case
We first benchmark MIGHTY against GCOPTER (which uses MINCO [wang2022geometrically]) on a simple corner-avoidance task with fixed start and goal (Fig. 1). For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [liu1989lbfgs]) and stopping tolerances. Both methods start from the same initial guess.
GCOPTER represents each segment as a fifth-order polynomial, so MINCO minimizes jerk (third-order effort) by construction. MIGHTY uses a fifth-order Hermite spline to match the polynomial degree. GCOPTER applies diffeomorphic parameterizations for waypoints and durations (waypoints remain inside the overlap of safe-flight corridors; durations stay positive). The only difference is that MIGHTY includes an explicit jerk smoothness term (Sec. III-C1), while GCOPTER does not add a separate smoothness cost because MINCO already minimizes jerk.
Figure 1: Benchmarking environment for a simple corner avoidance scenario. The blue-shaded polygons show the safe flight corridor. MIGHTY’s trajectory is colored by speed (warmer colors indicate faster speeds).
Figure 2: Benchmarking results for MIGHTY and GCOPTER for simple corner avoidance scenario in Fig 1. MIGHTY achieves lower computation time, travel time, and path length, while GCOPTER yields lower jerk.
[…]
The result summary in Fig. 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER. While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO manifold, MIGHTY explores a larger search space to identify faster and more effective trajectories. This flexibility enables MIGHTY to achieve superior overall performance.
IV-B Representation Benchmarking: Complex Case
Figure 3: Complex-scene benchmarking setup. MIGHTY’s trajectory is color-mapped by speed (warm=fast), and GCOPTER’s trajectory is blue. The blue-shaded polygons show the safe flight corridor. The start is at $(0,0,0.5)\,$\mathrm{m}$\$, and 24 goals lie on a grid with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=$2.5\text{\,}\mathrm{m}$\$. Blue-shaded polygons show the safe flight corridor sequence from one run.
TABLE II: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
$v_{\max}$ [m/s] Algo. $T_{\mathrm{comp}}$ [ms] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $\int\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2] $\rho_{\mathrm{viol}}$ [%]
1.0 GC 7.7 25.1 19.1 2.1 0.0
MI
8.1
(+4.8%)
21.5
(-14.2%)
18.8
(-1.4%)
6.9
(+229.0%)
0.0
(0.0%)
2.0 GC 9.8 12.6 19.0 7.9 0.0
MI
9.3
(-4.7%)
11.0
(-12.8%)
18.6
(-2.2%)
17.2
(+118.5%)
0.0
(0.0%)
3.0 GC 11.8 8.5 19.0 13.9 0.0
MI
7.8
(-33.6%)
7.7
(-9.8%)
18.7
(-1.4%)
27.2
(+95.9%)
0.0
(0.0%)
4.0 GC 7.7 6.9 18.8 17.6 0.0
MI
7.2
(-5.5%)
6.0
(-13.0%)
18.6
(-1.2%)
35.6
(+102.6%)
0.0
(0.0%)
5.0 GC 6.8 6.1 18.8 19.8 0.0
MI
7.2
(+5.7%)
5.3
(-14.1%)
18.7
(-0.7%)
43.7
(+120.5%)
0.0
(0.0%)
Overall -9.3% -13.1% -1.4% +113.3% 0.0%
We further evaluate MIGHTY in a complex environment (Fig. 3). The start is fixed at $(0,0,0.5)\,$\mathrm{m}$\$. For each $v_{\max}\in\{1,2,3,4,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, we plan to 24 goal locations with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=2.5\,$\mathrm{m}$\$. Optimization settings are identical to the simple case for both methods. We report the following metrics: $v_{\max}$ [m/s]: Speed limit, $T_{\mathrm{comp}}$ [ms]: Computation time, $T_{\mathrm{trav}}$ [s]: Travel time, $L_{\mathrm{path}}$ [m]: Path length, $\int\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2]: Jerk smoothness integral, and $\rho_{\mathrm{viol}}$ [%]: Fraction of time violating dynamic bounds.
[…]
IV-C Ablation Study on Scaled and Unscaled Variables
Figure 4: Ablation study comparing MIGHTY with scaled and unscaled variables in the complex benchmarking scenario. The scaled version shows $\approx 2\times$ faster computation time, with lower jerk and slightly shorter path length, while travel time is nearly identical.
As discussed in Section III-D, we optimize over scaled variables to improve numerical stability. We evaluate this with an ablation between MIGHTY (scaled) and an unscaled variant. As in the complex benchmarking, both run in the same environment with 24 goals for $v_{\max}\in\{1,\dots,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, using identical objectives, constraints, initial guess, and solver settings. Fig. 4 shows that the scaled variant is about $2\times$ faster in computation time, with lower jerk, while path length and travel time are nearly identical. These results confirm that scaling improves numerical stability and optimization efficiency without degrading trajectory quality.
IV-D Benchmarking in Static Environments
Figure 5: (Top) Static obstacle environment used for benchmarking. (Bottom) Top view of the point cloud and path generated by MIGHTY in the static environment. The warmer colors indicate faster speeds.
TABLE III: Benchmarking results (safety, computation time, performance, and constraint violation). We mark in green the best value in each column considering only configurations with $R_{\mathrm{succ}}=100\%$; configurations with $R_{\mathrm{succ}}<100\%$ are excluded from best highlighting. For EGO-Swarm2, we evaluated both its default sensor, depth camera, and a LiDAR (as used by SUPER and MIGHTY) and report results as “camera”  |  “LiDAR”. We also tested EGO-Swarm with LiDAR, but its performance was substantially worse, so we omit it from the table for clarity. For SUPER, computation time is reported as “exploratory”  &  “safe”. The data are shaded red if they have unsuccessful rates $R_{\mathrm{succ}}<100\%$.
Algorithm Params Traj. Success Computation Time Performance Constraint Violation
$R_{\mathrm{succ}}$ [%] $T_{\mathrm{opt}}$ [ms] $T_{\mathrm{total}}$ [s] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $S_{\mathrm{jerk}}$ [m/s2] $\bar{S}_{\mathrm{jerk}}$ [m/s3] $\rho_{\mathrm{vel}}$ [%] $\rho_{\mathrm{acc}}$ [%] $\rho_{\mathrm{jerk}}$ [%]
EGO-Swarm $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ B-spline 40.0 1.8 2.3 135.9 379.7 4233.1 43.6 3.5 6.5 38.0
$v_{\max}=$3.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ 70.0 1.8 2.4 178.2 380.2 3024.5 26.7 5.7 1.2 16.5
$v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ 80.0 0.9 1.3 262.7 364.8 1074.9 7.3 3.8 0.01 0.4
$v_{\max}=$1.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ 90.0 0.6 0.9 559.8 370.0 204.4 0.6 0.1 0.0 0.0
EGO-Swarm2 $w_{\rm obst}=$1\text{\times}{10}^{4}$\$ MINCO 100.0 | 20.0 1.3 | 5.9 1.8 | 7.6 144.7 | 144.6 329.8 | 331.7 263.3 | 336.8 2.5 | 3.5 0.0 | 0.6 0.0 | 0.0 0.0 | 0.01
$w_{\rm obst}=$1\text{\times}{10}^{5}$\$ 100.0 | 70.0 1.2 | 4.6 1.7 | 5.9 140.2 | 150.1 326.9 | 333.5 243.2 | 247.7 2.2 | 3.5 0.05 | 0.02 0.0 | 0.0 0.0 | 0.0
SUPER $w_{t}=$1\text{\times}{10}^{4}$\$ MINCO 20.0 5.9 & 18.0 46.4 87.5 323.5 627.4 10.9 0.0 0.02 1.0
$w_{t}=$1\text{\times}{10}^{3}$\$ 60.0 5.6 & 18.0 57.5 99.6 323.4 431.9 6.7 0.0 0.02 0.2
$w_{t}=$1\text{\times}{10}^{2}$\$ 60.0 6.8 & 18.2 47.6 124.7 326.9 399.5 4.2 0.0 0.01 0.06
$w_{t}=$1\text{\times}{10}^{1}$\$ 100.0 7.8 & 18.9 50.6 264.7 342.0 200.0 0.9 0.0 0.0 0.0
MIGHTY Hermite 100.0 10.5 19.7 79.0 310.9 522.2 9.4 0.0 0.0 0.0
To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig. 5) against EGO-Swarm [zhou2021ego-swarm], EGO-Swarm2 [zhou2022swarm], and SUPER [ren2025super]. Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area. The agent starts at $(0,0,3)\,$\mathrm{m}$\$ and the goal is $(305,0,3)\,$\mathrm{m}$\$. We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$. For global guidance we run JPS [harabor2011online], then build a safe-flight corridor (SFC) [liu2017planning] for local optimization.
For soft constraints we use a smooth hinge $\phi_{\mu}(\cdot)$. Time integrals are evaluated by trapezoidal quadrature with $\kappa$ samples per segment unless a closed form is available. Safe-flight corridors (SFCs) are modeled per segment $s$ as the intersection of halfspaces
$\text{SFC}_{s}\;\triangleq\;\big\{\mathbf{x}\in\mathbb{R}^{3}\;\big|\;\mathbf{a}_{s,h}^{\top}\mathbf{x}\leq b_{s,h}-C_{\text{SFC}},\ \ h=1,\dots,H_{s}\big\},$
where $H_{s}$ is the number of halfspaces on segment $s$, and $C_{\text{SFC}}=0.2$ is a safety margin. The total objective is
$J=$ $~w_{T}\!\sum_{s}T_{s}+w_{\text{smooth}}\tilde{J}_{\text{smooth}}+w_{\text{SFC}}J_{\text{SFC}}$
$~+w_{v}J_{v}+w_{a}J_{a}+w_{j}J_{j},$
where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{a}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$, Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq. 6, which avoids per-sample recomputation. MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\text{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$. We run 10 trials per method. Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s3] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations). We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.
[…]
IV-E Dynamic Environments
Figure 6: Benchmarking scenario with 100 dynamic obstacles following smooth trefoil-knot trajectories. The image shows the agent’s trajectory (color-mapped by speed) navigating through the dynamic obstacles.
Figure 7: Benchmarking scenario with 50 dynamic obstacles and static obstacles. The top left image illustrates the Gazebo simulation environment. The main image and bottom right image show the point cloud (pink-orange) and agent’s trajectory (color-mapped by speed) navigating through both static and dynamic obstacles.
TABLE IV: Nearest dynamic-obstacle distances by environment for 10 runs each. Columns show 0th (min), 5th, 50th (median), and 95th percentiles. The robot collision radius is set to $0.1\text{\,}\mathrm{m}$.
Environment Distance [m]
0th (min) 5th 50th (median) 95th
Only Dynamic 0.8 2.1 3.2 5.2
Dynamic + Static 1.0 2.5 3.8 6.8
We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles. For both, in addition to the cost terms described in Section IV-D, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$. We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$. Note that $C_{\text{dyn}}$ is where the penalty activates, not a hard safety distance. The robot collision radius is $0.1\text{\,}\mathrm{m}$. The start and goal are $(0,0,3)\,$\mathrm{m}$\$ and $(100,0,3)\,$\mathrm{m}$\$, and, in this more complex environment, the dynamic constraints are set to $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
Dynamic-only environment: We generate 100 dynamic obstacles following smooth trefoil-knot trajectories, with initial positions $x\in[0,100]\,$\mathrm{m}$\$, $y\in[-10,10]\,$\mathrm{m}$\$, $z\in[1,6]\,$\mathrm{m}$\$. Fig. 6 shows an example of the environment and the trajectory. MIGHTY avoids all obstacles in 10/10 trials. As shown in Table IV, the minimum nearest-obstacle distance is $0.8\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ safety threshold.
Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig. 7 and run 10 trials. MIGHTY avoids all obstacles in all trials. Table IV summarizes nearest-obstacle distances, where the minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
V Hardware Experiments
We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission. The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [chen2023dlio]. Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [meier2015px4] on a Pixhawk flight controller [meier2011pixhawk]. All perception, planning, and control run in real time onboard.
Figure 8: Long-duration experiment with $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$. Top: partial action sequence at $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ (cropped for clarity). Bottom: full trajectory histories overlaid on the recorded LiDAR point cloud. Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
Long-Duration Flight Experiment: To assess reliability over extended operations, the vehicle repeatedly traverses a space with multiple obstacles (Fig. 8). Six goal positions are placed opposite the start, inducing repeated out-and-back motions through the environment. For $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, all flights completed without collision.
Figure 9: High-speed experiments at $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Each panel shows the action sequence with trajectory history over the recorded point cloud. Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
High-Speed Flight Experiment: We further evaluate performance at higher speeds. As shown in Section IV, MIGHTY tends to produce higher-jerk trajectories; when constraints are satisfied, these trajectories remain safe while enabling faster motion. We test $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Figure 9 shows the resulting histories. All flights completed without collision, demonstrating that the MIGHTY planner and DLIO state estimation operate effectively at high speed. At $v_{\max}=$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, the vehicle reached $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ peak speed.
Figure 10: Dynamic-environment experiment at $v_{\max}=$1.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. The red circle shows the UAV; obstacles are introduced and moved during flight.
Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig. 10). We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-E for details. The agent successfully avoids all obstacles and reaches the goal without collision over $490\text{\,}\mathrm{s}$ of flight time.
VI Conclusions
We presented MIGHTY, a Hermite-spline trajectory planner for joint spatiotemporal optimization. By optimizing path geometry and time together with local derivative control, MIGHTY generates smooth, collision-free, and dynamically feasible trajectories for high-performance maneuvers. Benchmarks and hardware tests show MIGHTY outperforms baselines in travel time while satisfying constraints.
VII Acknowledgements
The authors would like to thank Lili Sun for her help with the hardware experiments and Mason B. Peterson for his support with robotdatapy library. The authors also thank Juan Rached for his help with hardware setup. This work was supported in part by DSTA and ARL grant W911NF-21-2-0150.

### v3

@@
 I Introduction
-Trajectory planning for autonomous navigation has been extensively studied, with a wide variety of parameterizations and formulations [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus, zhou2019robust, zhou2021ego-planner, zhou2022swarm, zhou2021raptor, tordesillas2022faster, richter2016polynomial, levy2024stitcher, wang2022geometrically, ren2025super].
-For real-world deployment, fast trajectory generation is critical to ensure agile and safe flight in cluttered and dynamic environments.
-Approaches that use hard constraints [mellinger2011minimum, 7138978, liu2017planning, 9147300, kondo2025dynus] explicitly enforce safety and physical feasibility but often rely on computationally intensive solvers, making them difficult to deploy for high-frequency replanning on onboard processors.
-Soft-constraint planners such as EGO-Planner [zhou2021ego-planner], RAPTOR [zhou2021raptor], and SUPER [ren2025super] achieve faster convergence and lower computational cost.
-Some jointly optimize path geometry and timing within a single nonlinear formulation [wang2022geometrically, ren2025super], while others separate path generation and time allocation into two stages [richter2016polynomial, zhou2021raptor].
-In general, increasing the number of decision variables, such as control points, segment durations, or derivative values, can improve performance but also enlarges the optimization problem and increases computation time.
-To balance efficiency and performance, recent work has explored alternative trajectory representations and optimization strategies, and building on these ideas, the following introduces our Hermite spline–based planner, MIGHTY.
+Trajectory planning for autonomous navigation has been extensively studied with various parameterizations [13, 3, 10, 18, 7, 23, 25, 26, 24, 19, 17, 8, 22, 16].
+Hard-constraint approaches [13, 3, 10, 18, 7] explicitly enforce safety but require computationally intensive solvers unsuitable for high-frequency replanning.
+Soft-constraint planners (EGO-Planner [25], RAPTOR [24], SUPER [16]) achieve faster convergence.
+Some jointly optimize geometry and timing [22, 16], while others decouple path and time allocation [17, 24].
+Increasing decision variables improves performance but enlarges the problem.
+Building on these ideas, we introduce MIGHTY, a Hermite spline-based planner.
 TABLE I: State-of-the-art Unconstrained UAV Trajectory Planners
 Method Representation Variables Time Opt. Search Space Local Control
-Polynomial [richter2016polynomial] (2016) 9th-order polynomial endpoint derivs, time decoupled limited polynomial+time space yes
-Robust [zhou2019robust] (2019) non-uniform cubic B-spline B-spline cntrl pts, time decoupled limited polynomial+time space yes
-EGO-Planner [zhou2021ego-planner] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
-RAPTOR [zhou2021raptor] (2021) uniform cubic B-spline B-spline cntrl pts fixed limited polynomial+time space yes
-MINCO class [wang2022geometrically, zhou2022swarm, ren2025super] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO manifold) no
-MIGHTY (proposed) Hermite Hermite cntrl pts, time. joint full polynomial+time space yes
+Geom. Deriv.
+Polynomial [17] (2016) 9th-order polynomial endpoint derivs, time decoupled limited polynomial+time space no (global) direct but global
+Robust [23] (2019) non-uniform B-spline B-spline cntrl pts, time decoupled limited polynomial+time space yes indirect
+EGO-Planner [25] (2021) uniform B-spline B-spline cntrl pts fixed limited polynomial+time space yes indirect
+RAPTOR [24] (2021) uniform B-spline B-spline cntrl pts fixed limited polynomial+time space yes indirect
+MINCO class [22, 26, 16] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO class) no (global) indirect
+MIGHTY (proposed) Hermite spline Hermite cntrl pts, time. joint full polynomial+time space yes direct & local
 I-A Review of Unconstrained Frameworks
-Differential flatness [FLIESS01061995] enables efficient optimization of quadrotor trajectories by parameterizing the flat output ($x$, $y$, $z$, yaw) and their derivatives.
-Given fixed timestamps at trajectory waypoints, minimum control effort (e.g., jerk/snap) problems using piecewise polynomial representation can be formulated as quadratic programs (QPs), which are convex and can be solved efficiently [mellinger2011minimum].
+Differential flatness [5] enables efficient optimization of quadrotor trajectories by parameterizing the flat output and their derivatives.
+Given fixed timestamps at trajectory waypoints, minimum control effort (e.g., jerk/snap) problems using piecewise polynomial representation can be formulated as quadratic programs (QPs), which are convex and can be solved efficiently [13].
 However, when segment durations are included as optimization variables (time allocation), the resulting joint problem becomes substantially more complex, yielding a nonlinear program (NLP) that is often ill-conditioned because of the coupling between spatial and temporal variables.
-One common strategy is to decouple the problem into two stages and relax hard constraints to adopt unconstrained formulations.
-For example, Richter et al.
-[richter2016polynomial] apply piecewise 9th-order polynomials on the flat outputs with endpoint derivatives as decision variables.
-The trajectory optimization follows a two-stage process: an unconstrained QP minimizes snap to ensure smoothness, followed by a gradient-based time allocation step that adjusts the total trajectory duration.
-Collision avoidance is addressed heuristically by inserting mid-segment waypoints whenever a segment intersects obstacles, which triggers repeated QP solves until a collision-free solution is found.
-This approach is extended in [7353622] by improving numerical stability and incorporating additional soft state constraint costs to further ensure trajectory feasibility.
-Similarly, the work in [7759784] adapts this framework to achieve local real-time replanning, using endpoint derivatives as optimization variables, which minimizes snap and collisions using a potential cost function based on Euclidean signed distance field (ESDF).
+Decoupling into two stages relaxes hard constraints for unconstrained formulations.
+Richter et al.
+[17] use 9th-order polynomials with endpoint derivatives: a QP minimizes snap, then gradient-based time allocation adjusts duration.
+Endpoint derivatives are decision variables, so there is direct derivative parameterization.
+However, the free derivatives are globally coupled, so there is no local control.
+Collision avoidance inserts mid-segment waypoints, triggering repeated solves.
+Extensions [1, 14] improve stability and enable real-time replanning via ESDF-based potential costs.
 Another standard unconstrained framework uses B-splines for trajectory representation, with time-related variables encoded in a knot vector.
-B-splines offer inherent smoothness, local control, and the convex hull property, which enables efficient collision checking and enforces dynamic feasibility [9147300, zhou2019robust, zhou2021ego-planner, zhou2021raptor].
+B-splines offer inherent smoothness and geometric local control: changing one control point affects only a bounded neighborhood of the curve.
+The convex hull property enables efficient collision checking and enforces dynamic feasibility [18, 23, 25, 24].
+However, B-splines do not provide direct local control of higher-order derivatives.
+Derivatives at any curve point are implicit functions of multiple neighboring control points.
+One can bound derivatives efficiently via the convex hull property, but cannot independently prescribe a specific velocity or acceleration at a given knot through a single variable.
+This geometric local control property has been widely exploited for efficient trajectory refinement and replanning in cluttered environments [6, 15, 20, 4].
 Zhou et al.
-[zhou2019robust] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths.
+[23] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths.
 The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration.
 Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility.
-EGO-Planner [zhou2021ego-planner] and RAPTOR [zhou2021raptor] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points.
-EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [liu1989lbfgs] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness.
+EGO-Planner [25] and RAPTOR [24] simplify the representation by using uniform B-splines with fixed knot spans, leveraging convex hull properties to enforce velocity, acceleration, and jerk constraints directly on control points.
+EGO-Planner optimizes trajectories via a quasi-Newton solver (L-BFGS) [9] for fast gradient-based convergence, with an additional time-reallocation and anisotropic curve-fitting refinement stage that resizes knot spans and fits a new spline to ensure feasibility and smoothness.
 RAPTOR also employs a two-stage process: a closed-form QP warm-start for initial smoothness, followed by a nonlinear refinement phase that penalizes collision via an ESDF and ensures dynamic limits via control point constraints.
-To further improve efficiency, alternative representations of polynomial variables have been explored.
 Wang et al.
-[wang2021glst] employ an explicit diffeomorphism between coefficients and boundary derivatives, with a linear-complexity scheme for gradient evaluation towards coefficients.
-Built on this, MINCO [wang2022geometrically] parametrizes a segmented trajectory as piecewise polynomials, where coefficients are analytically determined by intermediate waypoints and segment durations, which are jointly optimized.
-A linear-complexity solver then constructs the minimum-control trajectory that minimizes the integrated control effort (e.g., jerk/snap) subject to flatness boundary conditions.
-Spatiotemporal deformation leverages smooth diffeomorphisms and convex constraint-elimination maps to enforce arbitrary collision and state-input constraints.
-However, although MINCO provides an elegant and computationally efficient formulation, since it directly maps intermediate waypoints and segment durations to polynomial coefficients using the optimality conditions of the optimal control-effort problem [wang2022geometrically], the search is, by construction, restricted to a low-dimensional manifold, potentially limiting flexibility when additional constraints or objectives are introduced.
-Additionally, because MINCO determines all polynomial coefficients by solving a single banded linear system subject to high-order continuity and optimality conditions, any change to one intermediate waypoint or segment duration generally changes the entire solution.
-In other words, MINCO does not have strict local control, meaning a small change to one waypoint or duration can propagate to many segments, complicating per-segment constraint handling for local changes.
-EGO-Swarm2 [zhou2022swarm] and SUPER [ren2025super] both adopt MINCO to compactly parameterize trajectories, with SUPER further optimizing exploratory and backup trajectories in each replanning cycle to balance aggressiveness and guaranteed safety.
-However, the intrinsic coupling within the MINCO formulation remains a limitation.
-MIGHTY aims to address this by optimizing over a higher-dimensional representation that preserves local control and improves flexibility for handling complex constraints.
+[21] employ diffeomorphism between coefficients and derivatives with linear-complexity gradients.
+MINCO [22] builds on this: coefficients are analytically determined by waypoints and durations, jointly optimized via linear-complexity solver for minimum jerk/snap.
+MINCO excels when minimum control effort dominates, but searches only the MINCO class.
+MINCO’s waypoint positions are decision variables, but polynomial coefficients are determined through a banded linear system that couples adjacent segments, so there is no geometric local control.
+Higher-order derivatives are not decision variables, and they are analytically determined from waypoints and durations, so there is no direct higher-derivative control.
+Under added objectives, this restriction can yield suboptimal performance compared to more general representations.
+EGO-Swarm2 [26] and SUPER [16] adopt MINCO; SUPER adds exploratory/backup trajectories for safety.
+MIGHTY optimizes positions and derivatives at each knot and time durations as explicit variables, enabling local control of waypoints and higher-order dynamics.
 I-B MIGHTY Contributions
-Summarizing the discussion above, direct unconstrained higher-order polynomial representations [richter2016polynomial] are typically too slow for real-time planning.
-Uniform or Non-uniform splines [zhou2019robust, zhou2021ego-planner, zhou2021raptor] can generate trajectories faster compared to spatiotemporal joint optimization; however, they can generate sub-optimal trajectories due to decoupled time adjustment.
-Waypoint-and-duration-based parameterizations [wang2022geometrically, zhou2022swarm, ren2025super] are restricted to a low-dimensional manifold, and a small change in one waypoint propagates globally through all polynomial coefficients.
-This is undesirable for real-time planning because it potentially complicates constraint/cost handling and slows down convergence.
+Table I summarizes the discussion above.
+Unconstrained higher-order polynomial representations [17] are typically too slow for real-time planning.
+Uniform or Non-uniform splines [23, 25, 24] can generate trajectories faster compared to spatiotemporal joint optimization; however, they can generate sub-optimal trajectories due to decoupled time adjustment.
+Waypoint-and-duration-based parameterizations [22, 26, 16] are restricted to the low-dimensional subspace induced by the MINCO parameterization, lack direct parameterization of higher-order derivatives at individual knots, and exhibit waypoint propagation where changes to one waypoint affect multiple segments through the globally coupled coefficient system.
+While these effects enable compact representation and analytical optimality for minimum-control objectives, they can complicate constraint/cost handling beyond minimum-control and limit local control of trajectory shape and higher-order dynamics.
 To address the limitations of existing soft-constraint planners, MIGHTY uses a Hermite-spline representation and an unconstrained nonlinear optimizer that searches directly over spatial waypoints, endpoint derivatives, and segment durations.
-A Hermite spline representation also guarantees continuity without global coefficient coupling.
-As such, the key contributions of the paper are:
+A Hermite spline representation guarantees continuity without global coefficient coupling, which fundamentally enables direct local control over both positions and higher-order dynamics at each knot.
+These properties could be exploited for efficient trajectory refinement (as in B-spline methods [6, 15, 20, 4]) but are not available in MINCO’s globally coupled parameterization.
+The key contributions of the paper are:
   1. 1.
-    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, fully leveraging the continuous search space of Hermite splines and supporting local control of trajectory shape.
+    Hermite Spline Planner: Joint spatiotemporal optimization in a single solve, with explicit parameterization of positions and derivatives at each knot.
+This representation enables local control of higher-order dynamics and trajectory shape, which the optimizer can leverage when needed.
 The approach, MIGHTY, achieves fast solve times and superior performance.
@@
   3. 3.
-    Extensive Hardware Experiments: Long-duration flights, high-speed flights, and dynamic obstacle avoidance scenarios using a LiDAR-based perception and localization system.
-MIGHTY achieves collision-free long-duration flights and a top speed of $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ in high-speed flight experiments.
+    Extensive Hardware Experiments: Long-duration and high-speed flights, and dynamic obstacle avoidance scenarios using a LiDAR-based perception and localization system.
+MIGHTY achieves collision-free flights and a top speed of $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ in high-speed flight experiments.
 II Hermite Spline
@@
 III-C Objective and Closed-Form Gradient
-To make the formulation basis-independent, we express the objective and its gradients with respect to a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.).
+To make the formulation basis-independent, we express the objective and its gradients w.r.t a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.).
 As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
 $\nabla_{\mathbf{v}_{s+1}}J_{s}$ $=-\tfrac{2T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,3}}-\tfrac{T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,4}},\nabla_{\mathbf{a}_{s+1}}J_{s}=\tfrac{T_{s}^{2}}{20}\,\mathbf{g}_{\mathbf{c}_{s,3}}.$
-And $\partial\mathbf{c}_{s}/\partial T_{s}$ used in the coefficient path in Eq. (4) can be computed in closed form as
-$\frac{\mathbf{c}_{s,0}}{T_{s}}$ $=0,\quad\frac{\mathbf{c}_{s,1}}{T_{s}}=\tfrac{1}{5}\,\mathbf{v}_{s},\quad\frac{\mathbf{c}_{s,2}}{T_{s}}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
-$\frac{\mathbf{c}_{s,3}}{T_{s}}$ $=-\tfrac{2}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s+1},\quad\frac{\mathbf{c}_{s,4}}{T_{s}}=-\tfrac{1}{5}\,\mathbf{v}_{s+1},\quad\frac{\mathbf{c}_{s,5}}{T_{s}}=0.$
+And $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
+$\nabla_{T_{s}}\mathbf{c}_{s,0}$ $=\mathbf{0},\ \nabla_{T_{s}}\mathbf{c}_{s,1}=\tfrac{1}{5}\,\mathbf{v}_{s},\ \nabla_{T_{s}}\mathbf{c}_{s,2}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
+$\nabla_{T_{s}}\mathbf{c}_{s,3}$ $=-\tfrac{2}{5}\,\mathbf{v}_{s+1}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s+1},\ \nabla_{T_{s}}\mathbf{c}_{s,4}=-\tfrac{1}{5}\,\mathbf{v}_{s+1},$
+$\nabla_{T_{s}}\mathbf{c}_{s,5}$ $=\mathbf{0}.$
 Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Section IV.
@@
 $\bar{T}_{i}=\begin{cases}T_{0},&i=0,\\ \frac{T_{i-1}+T_{i}}{2},&i=1,\dots,M-1,\\ T_{M-1},&i=M,\end{cases}$
-and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization removes the explicit $1/T_{s}$ factors from the gradients and balances sensitivity across short and long segments.
-Our ablation study in Section IV-C confirms the expected improvement in performance.
+and set $\hat{\mathbf{v}}_{i}=\bar{T}_{i}\,\mathbf{v}_{i},\quad\hat{\mathbf{a}}_{i}=\bar{T}_{i}^{2}\,\mathbf{a}_{i}.$ This reparameterization aims to reduce the explicit $1/T_{s}$ and $1/T_{s}^{2}$ factors’ effect in the gradients w.r.t.
+the derivative decision variables by optimizing the scaled variables $\{\hat{\mathbf{v}}_{i},\hat{\mathbf{a}}_{i}\}$ instead of $\{\mathbf{v}_{i},\mathbf{a}_{i}\}$.
+At the endpoints, the scaling uses the adjacent segment duration (no averaging), while at interior knots we use the local average duration to balance the influence of the two neighboring segments; in that case, the inverse-duration effects are mitigated but cannot be canceled for both segments simultaneously unless their durations match.
+Our ablation study in Section IV-E confirms the expected improvement in performance.
 IV Simulation Results
@@
 IV-A Representation Benchmarking: Simple Case
-We first benchmark MIGHTY against GCOPTER (which uses MINCO [wang2022geometrically]) on a simple corner-avoidance task with fixed start and goal (Fig.
+We first benchmark MIGHTY against GCOPTER (which uses MINCO [22]) on a simple corner-avoidance task with fixed start and goal (Fig.
 1).
-For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [liu1989lbfgs]) and stopping tolerances.
+For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [9]) and stopping tolerances.
 Both methods start from the same initial guess.
@@
 […]
-The result summary in Fig.
+The result in Fig.
 2 shows that MIGHTY achieves a shorter travel time, with slightly faster computation and shorter path than GCOPTER.
-While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO manifold, MIGHTY explores a larger search space to identify faster and more effective trajectories.
-This flexibility enables MIGHTY to achieve superior overall performance.
-IV-B Representation Benchmarking: Complex Case
+While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO class, MIGHTY explores a larger search space to identify faster and more effective trajectories.
+This flexibility enables MIGHTY to achieve superior performance, and Section IV-B further investigates the jerk-performance trade-off.
+IV-B Smooth Weight Sweep Benchmarking
+Fig.
+2 shows MIGHTY yields higher jerk than GCOPTER while achieving better travel time.
+To quantify this trade-off, we sweep $w_{\mathrm{smooth}}\in\{${10}^{-5}$,\ldots,${10}^{3}$\$\} in the same scenario.
+Fig.
+3 shows that as $w_{\mathrm{smooth}}$ increases, MIGHTY progressively reduces jerk, eventually surpassing GCOPTER’s levels of smoothness.
+At $w_{\mathrm{smooth}}=${10}^{2}$\$, MIGHTY matches GCOPTER’s jerk while maintaining better performance and lower computation time, demonstrating that MIGHTY’s higher jerk is a tunable design choice rather than an inherent limitation.
+Figure 3: Effect of jerk-smoothness weighting in the simple corner-avoidance scenario (Fig.
+1).
+As $w_{\mathrm{smooth}}$ increases, MIGHTY’s jerk decreases toward GCOPTER’s jerk (orange dotted reference).
+At $w_{\mathrm{smooth}}=${10}^{2}$\$ (vertical red line), MIGHTY matches GCOPTER’s jerk with comparable performance and lower computation time.
+IV-C Local Control of Higher-Order Dynamics
+MINCO-based planners lack direct parameterization of higher-order derivatives at individual knots; these quantities are derived from a globally coupled coefficient system (Sec.
+I-A).
+While GCOPTER can add derivative constraints as soft penalties, MIGHTY’s Hermite parameterization treats knot positions and velocities as explicit optimization variables.
+To evaluate this difference, we extend the simple corner-avoidance scenario (Fig.
+1) with soft reference tracking terms for position and velocity at the penultimate knot, which is relevant for waypoint velocity specification and hover-at-goal behaviors.
+Figure 4: Local control of higher-order dynamics: position and velocity reference tracking.
+GCOPTER and MIGHTY trajectories are color-mapped by speed.
+Both planners enforce a velocity reference $\mathbf{v}_{\text{ref}}=(0,0,0)\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$ and position reference at the penultimate knot (one knot before the goal).
+As shown, MIGHTY achieves better reference tracking accuracy.
+We augment the objective function with soft position and velocity reference terms
+$J_{\text{ref}}=\tfrac{1}{2}w_{\mathbf{p}_{\text{ref}}}\lVert\mathbf{p}(t_{k})-\mathbf{p}_{\text{ref}}\rVert^{2}+\tfrac{1}{2}w_{\mathbf{v}_{\text{ref}}}\lVert\mathbf{v}(t_{k})-\mathbf{v}_{\text{ref}}\rVert^{2},$
+where $\mathbf{p}_{\text{ref}}$ and $\mathbf{v}_{\text{ref}}$ are desired position and velocity at knot $k$, and $w_{\mathbf{p}_{\text{ref}}}$ and $w_{\mathbf{v}_{\text{ref}}}$ are penalty weights.
+For this benchmark, we set $k$ to the penultimate knot (one knot before the goal), $\mathbf{v}_{\text{ref}}=(0,0,0)\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$ to enforce near-hover at the goal, and $\mathbf{p}_{\text{ref}}$ to the goal position.
+We use $w_{\mathbf{p}_{\text{ref}}}=$1\text{\times}{10}^{3}$\$ and $w_{\mathbf{v}_{\text{ref}}}=$1\text{\times}{10}^{3}$\$, with $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ and $w_{\mathrm{smooth}}=${10}^{-5}$\$ for MIGHTY.
+All other settings match those in Sec.
+IV-F.
+We evaluate both GCOPTER and MIGHTY across 100 runs in the same simple corner-avoidance environment.
+Fig.
+4 shows example trajectories from both planners, and Table II summarizes the results.
+TABLE II: Local control of higher-order dynamics benchmark. We added position/velocity soft penalties on both GCOPTER and MIGHTY and evaluated performance. Best results are in green.
+Planner $e_{\mathbf{p}_{\text{ref}}}$ [m] $e_{\mathbf{v}_{\text{ref}}}$ [m/s] $T_{\mathrm{trav}}$ [s] $T_{\mathrm{comp}}$ [ms]
+GCOPTER 0.017 1.0 19.7 7.4
+MIGHTY 0.0041 0.16 14.8 10.6
+[…]
+IV-D Representation Benchmarking: Complex Case
 Figure 5: Complex-scene benchmarking setup.
@@
 Blue-shaded polygons show the safe flight corridor sequence from one run.
-TABLE II: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
+TABLE III: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
 $v_{\max}$ [m/s] Algo. $T_{\mathrm{comp}}$ [ms] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $\int\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2] $\rho_{\mathrm{viol}}$ [%]
@@
 We further evaluate MIGHTY in a complex environment (Fig.
-3).
-The start is fixed at $(0,0,0.5)\,$\mathrm{m}$\$.
-For each $v_{\max}\in\{1,2,3,4,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, we plan to 24 goal locations with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=2.5\,$\mathrm{m}$\$.
+5) with start at $(0,0,0.5)\,$\mathrm{m}$\$ and 24 goals at $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=2.5\,$\mathrm{m}$\$.
 Optimization settings are identical to the simple case for both methods.
@@
 […]
-IV-C Ablation Study on Scaled and Unscaled Variables
+IV-E Ablation Study on Scaled and Unscaled Variables
 Figure 6: Ablation study comparing MIGHTY with scaled and unscaled variables in the complex benchmarking scenario.
@@
 As discussed in Section III-D, we optimize over scaled variables to improve numerical stability.
-We evaluate this with an ablation between MIGHTY (scaled) and an unscaled variant.
-As in the complex benchmarking, both run in the same environment with 24 goals for $v_{\max}\in\{1,\dots,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, using identical objectives, constraints, initial guess, and solver settings.
+We evaluate this via ablation using the same complex benchmarking setup.
 Fig.
-4 shows that the scaled variant is about $2\times$ faster in computation time, with lower jerk, while path length and travel time are nearly identical.
-These results confirm that scaling improves numerical stability and optimization efficiency without degrading trajectory quality.
-IV-D Benchmarking in Static Environments
+6 shows the scaled variant is $\approx 2\times$ faster with lower jerk, while path length and travel time are nearly identical, confirming that scaling improves efficiency without degrading quality.
+IV-F Benchmarking in Static Environments
 Figure 7: (Top) Static obstacle environment used for benchmarking.
@@
 The warmer colors indicate faster speeds.
-TABLE III: Benchmarking results (safety, computation time, performance, and constraint violation).
+TABLE IV: Benchmarking results (safety, computation time, performance, and constraint violation).
 We mark in green the best value in each column considering only configurations with $R_{\mathrm{succ}}=100\%$; configurations with $R_{\mathrm{succ}}<100\%$ are excluded from best highlighting.
@@
 To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig.
-5) against EGO-Swarm [zhou2021ego-swarm], EGO-Swarm2 [zhou2022swarm], and SUPER [ren2025super].
+7) against EGO-Swarm [27], EGO-Swarm2 [26], and SUPER [16].
 Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area.
@@
 We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
-For global guidance we run JPS [harabor2011online], then build a safe-flight corridor (SFC) [liu2017planning] for local optimization.
+For global planning we run A*, then build a safe-flight corridor (SFC) [10] for local optimization.
+The intermediate waypoints for A* are projected from the terminal goal onto the local occupancy map centered on the drone.
+The average computation time for A* and SFC generation is $0.26\text{\,}\mathrm{m}\mathrm{s}$ and $0.48\text{\,}\mathrm{m}\mathrm{s}$, respectively.
 For soft constraints we use a smooth hinge $\phi_{\mu}(\cdot)$.
@@
 where $H_{s}$ is the number of halfspaces on segment $s$, and $C_{\text{SFC}}=0.2$ is a safety margin. The total objective is
-$J=$ $~w_{T}\!\sum_{s}T_{s}+w_{\text{smooth}}\tilde{J}_{\text{smooth}}+w_{\text{SFC}}J_{\text{SFC}}$
+$J=$ $~w_{T}\!\sum_{s}T_{s}+w_{\mathrm{smooth}}\tilde{J}_{\text{smooth}}+w_{\text{SFC}}J_{\text{SFC}}$
 $~+w_{v}J_{v}+w_{a}J_{a}+w_{j}J_{j},$
@@
 6, which avoids per-sample recomputation.
-MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\text{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$.
+MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\mathrm{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$.
 We run 10 trials per method.
@@
 […]
-IV-E Dynamic Environments
+IV-G Dynamic Environments
 Figure 8: Benchmarking scenario with 100 dynamic obstacles following smooth trefoil-knot trajectories.
@@
 The main image and bottom right image show the point cloud (pink-orange) and agent’s trajectory (color-mapped by speed) navigating through both static and dynamic obstacles.
-TABLE IV: Nearest dynamic-obstacle distances by environment for 10 runs each.
-Columns show 0th (min), 5th, 50th (median), and 95th percentiles.
-The robot collision radius is set to $0.1\text{\,}\mathrm{m}$.
-Environment Distance [m]
-0th (min) 5th 50th (median) 95th
-Only Dynamic 0.8 2.1 3.2 5.2
-Dynamic + Static 1.0 2.5 3.8 6.8
 We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles.
-For both, in addition to the cost terms described in Section IV-D, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$.
+For both, in addition to the cost terms described in Section IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$.
 We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$.
@@
 MIGHTY avoids all obstacles in 10/10 trials.
-As shown in Table IV, the minimum nearest-obstacle distance is $0.8\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ safety threshold.
+The minimum nearest-obstacle distance is $0.8\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ safety threshold.
 Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig.
@@
 MIGHTY avoids all obstacles in all trials.
-Table IV summarizes nearest-obstacle distances, where the minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
+The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
 V Hardware Experiments
 We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission.
-The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [chen2023dlio].
-Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [meier2015px4] on a Pixhawk flight controller [meier2011pixhawk].
+The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [2].
+Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [11] on a Pixhawk flight controller [12].
 All perception, planning, and control run in real time onboard.
 Figure 10: Long-duration experiment with $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$.
-Top: partial action sequence at $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ (cropped for clarity).
-Bottom: full trajectory histories overlaid on the recorded LiDAR point cloud.
+Top: action sequence at $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$.
+Bottom: trajectory histories overlaid on the recorded LiDAR point cloud.
 Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
 Long-Duration Flight Experiment: To assess reliability over extended operations, the vehicle repeatedly traverses a space with multiple obstacles (Fig.
-8).
+10).
 Six goal positions are placed opposite the start, inducing repeated out-and-back motions through the environment.
 For $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, all flights completed without collision.
-Figure 9: High-speed experiments at $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$.
+Figure 11: High-speed experiments at $v_{\max}=$5.0\text{\,}$,$6.0\text{\,}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$.
 Each panel shows the action sequence with trajectory history over the recorded point cloud.
-Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
 High-Speed Flight Experiment: We further evaluate performance at higher speeds.
@@
 Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig.
-10).
-We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-E for details.
-The agent successfully avoids all obstacles and reaches the goal without collision over $490\text{\,}\mathrm{s}$ of flight time.
+12).
+We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-G for details.
+The agent successfully reaches the goal without collision over $490\text{\,}\mathrm{s}$.
 VI Conclusions
@@
 Benchmarks and hardware tests show MIGHTY outperforms baselines in travel time while satisfying constraints.
-VII Acknowledgements
-The authors would like to thank Lili Sun for her help with the hardware experiments and Mason B.
-Peterson for his support with robotdatapy library.
-The authors also thank Juan Rached for his help with hardware setup.
-This work was supported in part by DSTA and ARL grant W911NF-21-2-0150.

### v4

@@
 MINCO class [13, 8, 14] (2022, 2022, 2025) MINCO pos., time joint limited polynomial+time space (MINCO class) no (global) indirect
-MIGHTY (proposed) Hermite spline Hermite cntrl pts, time. joint full polynomial+time space yes direct & local
+MIGHTY (proposed) Hermite spline Hermite cntrl pts, time joint full polynomial+time space yes direct & local
 I-A Review of Unconstrained Frameworks
@@
 However, when segment durations are included as optimization variables (time allocation), the resulting joint problem becomes substantially more complex, yielding a nonlinear program (NLP) that is often ill-conditioned because of the coupling between spatial and temporal variables.
-Decoupling into two stages relaxes hard constraints for unconstrained formulations.
+Decoupling into two stages enables unconstrained formulations by replacing hard constraints with soft penalties.
 Richter et al.
@@
 [6] parameterize trajectories as non-uniform cubic B-splines initialized from kinodynamic A* paths.
-The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from an Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration.
+The nonlinear optimization minimizes the cost considering smoothness, a repulsive collision potential from a Euclidean distance field (EDF), and soft penalties on derivative control points exceeding maximum velocity and acceleration.
 Finally, an iterative time-adjustment method rescales knot spans of the non-uniform spline to guarantee dynamic feasibility.
@@
 This representation enables local control of higher-order dynamics and trajectory shape, which the optimizer can leverage when needed.
-The approach, MIGHTY, achieves fast solve times and superior performance.
+The approach, MIGHTY, achieves fast solve times and lower travel time than baselines.
   2. 2.
@@
 MIGHTY uses a degree-$d$ Hermite spline with $M$ segments.
-For odd degree $d=2k+1$, each segment is parameterized by the knot values of the position and its derivatives up to order $k$.
-For example, a quintic Hermite spline ($d=5$, $k=2$) specifies position, velocity, and acceleration.
-Across knots, a Hermite spline automatically guarantees $\mathcal{C}^{k}$ continuity by sharing the knot values.
+For odd degree $d=2\nu+1$, each segment is parameterized by the knot values of the position and its derivatives up to order $\nu$.
+For example, a quintic Hermite spline ($d=5$, $\nu=2$) specifies position, velocity, and acceleration.
+Across knots, a Hermite spline automatically guarantees $\mathcal{C}^{\nu}$ continuity by sharing the knot values.
 Although Hermite splines apply in any dimension, we work in $\mathbb{R}^{3}$ for quadrotor planning and use $d=5$ in this paper.
@@
 $\quad\mathbf{x}_{s}(\tau_{s})=\sum_{k=0}^{5}h_{k}(\tau_{s})\,\mathbf{H}_{s,k},$ (1)
-where $\mathbf{H}_{s,0}=\mathbf{p}_{s},\mathbf{H}_{s,1}=T_{s}\,\mathbf{v}_{s},\mathbf{H}_{s,2}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s},\mathbf{H}_{s,3}=\mathbf{p}_{s+1},\ \mathbf{H}_{s,4}=T_{s}\,\mathbf{v}_{s+1},\ \mathbf{H}_{s,5}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s+1}.$ where $h_{k}(\tau_{s})$ are the standard quintic Hermite basis functions.
+where $\mathbf{H}_{s,0}=\mathbf{p}_{s},\mathbf{H}_{s,1}=T_{s}\,\mathbf{v}_{s},\mathbf{H}_{s,2}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s},\mathbf{H}_{s,3}=\mathbf{p}_{s+1},\ \mathbf{H}_{s,4}=T_{s}\,\mathbf{v}_{s+1},\ \mathbf{H}_{s,5}=\tfrac{1}{2}T_{s}^{2}\,\mathbf{a}_{s+1}.$ Here, $h_{k}(\tau_{s})$ are the standard quintic Hermite basis functions.
 The $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}\in\mathbb{R}^{3}\;(i=0,\dots,M)$ and $T_{s}>0\;(s=0,\dots,M-1)$ are the interior knot positions, velocities, accelerations, and segment durations, respectively.
@@
 III-A Decision Variables
-As shown in Section II, the knot states $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}(i=0,\dots,M)$ and segment durations $T_{s}(s=0,\dots,M-1)$ fully determine the trajectory.
+As shown in Sec.
+II, the knot states $\mathbf{p}_{i},\mathbf{v}_{i},\mathbf{a}_{i}(i=0,\dots,M)$ and segment durations $T_{s}(s=0,\dots,M-1)$ fully determine the trajectory.
 Thus we optimize the interior positions, velocities, accelerations, and per-segment durations:
-$\mathbf{z}=\bigl[\,\mathbf{p}_{1}^{\top},\mathbf{v}_{1}^{\top},\mathbf{a}_{1}^{\top},\ldots,\mathbf{p}_{M-1}^{\top},\mathbf{v}_{M-1}^{\top},\mathbf{a}_{M-1}^{\top},T_{0},\ldots,T_{M-1}\bigr].^{\top}$
+$\mathbf{z}=\bigl[\,\mathbf{p}_{1}^{\top},\mathbf{v}_{1}^{\top},\mathbf{a}_{1}^{\top},\ldots,\mathbf{p}_{M-1}^{\top},\mathbf{v}_{M-1}^{\top},\mathbf{a}_{M-1}^{\top},T_{0},\ldots,T_{M-1}\bigr]^{\top}.$
 with fixed boundary states $(\mathbf{p}_{0},\mathbf{v}_{0},\mathbf{a}_{0})$ and $(\mathbf{p}_{M},\mathbf{v}_{M},\mathbf{a}_{M})$.
 To keep the optimization unconstrained while enforcing certain requirements (e.g., $T_{s}>0$), we apply diffeomorphism, such as $T_{s}\;=\;\phi(\sigma_{s})$ with $\phi(\sigma)=e^{\sigma},\ \sigma_{s}\in\mathbb{R}.$ We also find that optimizing raw derivative knots $(\mathbf{v}_{i},\mathbf{a}_{i})$ can be numerically unstable.
-Section III-D introduces scalings that improve stability without changing the optimum.
+Sec.
+III-D introduces scalings that improve stability without changing the optimum.
 For clarity, the formulas below are written in terms of the original variables $(\mathbf{p}_{s},\mathbf{v}_{s},\mathbf{a}_{s},T_{s})$; applying a diffeomorphism only inserts chain-rule factors in the gradients.
@@
 III-C Objective and Closed-Form Gradient
-To make the formulation basis-independent, we express the objective and its gradients w.r.t a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.).
+To make the formulation basis-independent, we express the objective and its gradients w.r.t.
+a generic control-point (or coefficient) vector $\mathbf{c}$, independent of the chosen parameterization (Hermite end states, Bézier control points, etc.).
 As a concrete example, the affine map between Hermite and Bézier on segment $s$ is
@@
 $\nabla_{\mathbf{v}_{s+1}}J_{s}$ $=-\tfrac{2T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,3}}-\tfrac{T_{s}}{5}\,\mathbf{g}_{\mathbf{c}_{s,4}},\nabla_{\mathbf{a}_{s+1}}J_{s}=\tfrac{T_{s}^{2}}{20}\,\mathbf{g}_{\mathbf{c}_{s,3}}.$
-And $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
+The coefficient gradient $\partial\mathbf{c}_{s}/\partial T_{s}$ ($\nabla_{T_{s}}\mathbf{c}_{s}$) used in the coefficient path in Eq. (4) can be computed in closed form as
 $\nabla_{T_{s}}\mathbf{c}_{s,0}$ $=\mathbf{0},\ \nabla_{T_{s}}\mathbf{c}_{s,1}=\tfrac{1}{5}\,\mathbf{v}_{s},\ \nabla_{T_{s}}\mathbf{c}_{s,2}=\tfrac{2}{5}\,\mathbf{v}_{s}+\tfrac{T_{s}}{10}\,\mathbf{a}_{s},$ (5)
@@
 $\nabla_{T_{s}}\mathbf{c}_{s,5}$ $=\mathbf{0}.$
-Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Section IV.
+Therefore, to obtain the gradient, we need to compute $\mathbf{g}_{\mathbf{c}_{s}}$ and $\bigl.\partial J_{s}/\partial T_{s}\bigr|_{\text{explicit}}.$ Here we present two types of cost terms: (1) cost terms defined on control points, and (2) cost terms defined via sampled states along the trajectory, and their gradients used in Sec.
+IV.
 III-C1 Cost on Control Points
@@
 where $C_{s}=3600\,T_{s}^{-5}$.
-Define the (scalar) Bernstein-quadratic Gram matrix $G\in\mathbb{R}^{3\times 3}$ by $G_{mn}\;\triangleq\;\int_{0}^{1}B_{m}^{2}(\tau_{s})\,B_{n}^{2}(\tau_{s})\,d\tau_{s}\quad(m,n=0,1,2),$ so that, with $\Delta_{s}\triangleq$ then $\Delta_{s,0}^{\top},\Delta_{s,1}^{\top},\Delta_{s,2}^{\top}]^{\top}\in\mathbb{R}^{9}$,
+Define the (scalar) Bernstein-quadratic Gram matrix $G\in\mathbb{R}^{3\times 3}$ by $G_{mn}\;\triangleq\;\int_{0}^{1}B_{m}^{2}(\tau_{s})\,B_{n}^{2}(\tau_{s})\,d\tau_{s}\quad(m,n=0,1,2),$ so that, with $\Delta_{s}\triangleq[\Delta_{s,0}^{\top},\Delta_{s,1}^{\top},\Delta_{s,2}^{\top}]^{\top}\in\mathbb{R}^{9}$,
 $J_{\text{smooth}}\;=\;\sum_{s=0}^{M-1}C_{s}\;\Delta_{s}^{\top}\,Q\,\Delta_{s},\qquad Q\;\triangleq\;G\otimes I_{3}.$ (6)
@@
 At the endpoints, the scaling uses the adjacent segment duration (no averaging), while at interior knots we use the local average duration to balance the influence of the two neighboring segments; in that case, the inverse-duration effects are mitigated but cannot be canceled for both segments simultaneously unless their durations match.
-Our ablation study in Section IV-E confirms the expected improvement in performance.
+Our ablation study in Sec.
+IV-E confirms the expected improvement in performance.
 IV Simulation Results
@@
 1).
-For a fair comparison, we implemented GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [9]) and stopping tolerances.
+For a fair comparison, we implement GCOPTER’s optimization cost in MIGHTY and use the same optimizer (L-BFGS [22]) and stopping tolerances.
 Both methods start from the same initial guess.
@@
 MIGHTY uses a fifth-order Hermite spline to match the polynomial degree.
-GCOPTER applies diffeomorphic parameterizations for waypoints and durations (waypoints remain inside the overlap of safe-flight corridors; durations stay positive).
+GCOPTER applies diffeomorphic parameterizations for waypoints and durations (waypoints remain inside the overlap of safe flight corridors (SFCs); durations stay positive).
 The only difference is that MIGHTY includes an explicit jerk smoothness term (Sec.
 III-C1), while GCOPTER does not add a separate smoothness cost because MINCO already minimizes jerk.
-Figure 1: Benchmarking environment for a simple corner avoidance scenario.
-The blue-shaded polygons show the safe flight corridor.
-MIGHTY’s trajectory is colored by speed (warmer colors indicate faster speeds).
-Figure 2: Benchmarking results for MIGHTY and GCOPTER for simple corner avoidance scenario in Fig 1.
+Figure 1: Benchmarking environment for a simple corner avoidance scenario. The blue-shaded polygons show the SFC. MIGHTY’s trajectory is colored by speed (warmer colors indicate faster speeds).
+Figure 2: Benchmarking results for MIGHTY and GCOPTER for simple corner avoidance scenario in Fig.
+1.
 MIGHTY achieves lower computation time, travel time, and path length, while GCOPTER yields lower jerk.
@@
 While GCOPTER produces smoother trajectories with lower jerk due to its constraint within the MINCO class, MIGHTY explores a larger search space to identify faster and more effective trajectories.
-This flexibility enables MIGHTY to achieve superior performance, and Section IV-B further investigates the jerk-performance trade-off.
+This flexibility enables MIGHTY to achieve superior performance, and Sec.
+IV-B further investigates the jerk-performance trade-off.
 IV-B Smooth Weight Sweep Benchmarking
@@
 MIGHTY’s trajectory is color-mapped by speed (warm=fast), and GCOPTER’s trajectory is blue.
-The blue-shaded polygons show the safe flight corridor.
+The blue-shaded polygons show the SFC.
 The start is at $(0,0,0.5)\,$\mathrm{m}$\$, and 24 goals lie on a grid with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=$2.5\text{\,}\mathrm{m}$\$.
-Blue-shaded polygons show the safe flight corridor sequence from one run.
 TABLE III: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
@@
 The scaled version shows $\approx 2\times$ faster computation time, with lower jerk and slightly shorter path length, while travel time is nearly identical.
-As discussed in Section III-D, we optimize over scaled variables to improve numerical stability.
+As discussed in Sec.
+III-D, we optimize over scaled variables to improve numerical stability.
 We evaluate this via ablation using the same complex benchmarking setup.
@@
 Algorithm Params Traj. Success Computation Time Performance Constraint Violation
-$R_{\mathrm{succ}}$ [%] $T_{\mathrm{opt}}$ [ms] $T_{\mathrm{total}}$ [s] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $S_{\mathrm{jerk}}$ [m/s2] $\bar{S}_{\mathrm{jerk}}$ [m/s3] $\rho_{\mathrm{vel}}$ [%] $\rho_{\mathrm{acc}}$ [%] $\rho_{\mathrm{jerk}}$ [%]
+$R_{\mathrm{succ}}$ [%] $T_{\mathrm{opt}}$ [ms] $T_{\mathrm{total}}$ [ms] $T_{\mathrm{trav}}$ [s] $L_{\mathrm{path}}$ [m] $S_{\mathrm{jerk}}$ [m/s2] $\bar{S}_{\mathrm{jerk}}$ [m/s3] $\rho_{\mathrm{vel}}$ [%] $\rho_{\mathrm{acc}}$ [%] $\rho_{\mathrm{jerk}}$ [%]
 EGO-Swarm $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ B-spline 40.0 1.8 2.3 135.9 379.7 4233.1 43.6 3.5 6.5 38.0
@@
 We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
-For global planning we run A*, then build a safe-flight corridor (SFC) [10] for local optimization.
+For global planning we run A*, then build an SFC [3] for local optimization.
 The intermediate waypoints for A* are projected from the terminal goal onto the local occupancy map centered on the drone.
@@
 Time integrals are evaluated by trapezoidal quadrature with $\kappa$ samples per segment unless a closed form is available.
-Safe-flight corridors (SFCs) are modeled per segment $s$ as the intersection of halfspaces
-$\text{SFC}_{s}\;\triangleq\;\big\{\mathbf{x}\in\mathbb{R}^{3}\;\big|\;\mathbf{a}_{s,h}^{\top}\mathbf{x}\leq b_{s,h}-C_{\text{SFC}},\ \ h=1,\dots,H_{s}\big\},$
+SFCs are modeled per segment $s$ as the intersection of halfspaces
+$\text{SFC}_{s}\;\triangleq\;\big\{\mathbf{x}\in\mathbb{R}^{3}\;\big|\;\mathbf{n}_{s,h}^{\top}\mathbf{x}\leq b_{s,h}-C_{\text{SFC}},\ \ h=1,\dots,H_{s}\big\},$
 where $H_{s}$ is the number of halfspaces on segment $s$, and $C_{\text{SFC}}=0.2$ is a safety margin. The total objective is
@@
 $~+w_{v}J_{v}+w_{a}J_{a}+w_{j}J_{j},$
-where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{a}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$, Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq.
-6, which avoids per-sample recomputation.
+where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{n}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$.
+Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq.
+(6), which avoids per-sample recomputation.
 MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\mathrm{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$.
 We run 10 trials per method.
-Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s3] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations).
+Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations).
 We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.
@@
 We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles.
-For both, in addition to the cost terms described in Section IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$.
+For both, in addition to the cost terms described in Sec.
+IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$.
 We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$.
@@
 MIGHTY avoids all obstacles in all trials.
-The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
+The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ collision radius.
 V Hardware Experiments
-We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission.
+We also evaluated MIGHTY in three hardware experiments: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission.
 The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [25].
@@
 High-Speed Flight Experiment: We further evaluate performance at higher speeds.
-As shown in Section IV, MIGHTY tends to produce higher-jerk trajectories; when constraints are satisfied, these trajectories remain safe while enabling faster motion.
+As shown in Sec.
+IV, MIGHTY tends to produce higher-jerk trajectories; when constraints are satisfied, these trajectories remain safe while enabling faster motion.
 We test $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$.
-Figure 11 shows the resulting histories.
+Fig.
+11 shows the resulting histories.
 All flights completed without collision, demonstrating that the MIGHTY planner and DLIO state estimation operate effectively at high speed.
@@
 12).
-We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-G for details.
+We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Sec.
+IV-G for details.
 The agent successfully reaches the goal without collision over $490\text{\,}\mathrm{s}$.
@@
 Benchmarks and hardware tests show MIGHTY outperforms baselines in travel time while satisfying constraints.
+VII Acknowledgements
+The authors would like to thank Lili Sun for her help with the hardware experiments and Mason B.
+Peterson for his support with robotdatapy library.
+The authors also thank Juan Rached for his help with hardware setup.
+This work was supported in part by DSTA and ARL grant W911NF-21-2-0150.